the length of a rectangle is two more than three times its width the perimeter is 188. what is the length?
Answer by mananth(16949)


Let width be x
the length will be 3x+2
2x+2*(3x+2)=188
2x+6x+4=188
8x=188-4
x=184/8
x=23 whixh is the width
length = 71